In 2016, Vaughn Manufacturing sold 3000 units at $1000 each. Variable expenses were $560 per unit, and fixed expenses were $780000. The same variable expenses per unit and fixed expenses are expected for 2017. If Vaughn cuts selling price by 4%, what is Vaughn’s break-even point in units for 2017?

a. 1773.

b. 1875.

c. 1950.

d. 1847.
Answer #1

New selling price=1000*(1-0.04)=$960

Contribution margin=Sales-Variable cost

=960-560=$400 per unit

Breakeven=Fixed expenses/Contribution margin

=780000/400

=1950 units